# Condensing and Extracting Against Online Adversaries

### v2

@@ 1.2.3 Transforming Low-Entropy oNOSF sources to uniform oNOSF sources
Previously, [CGR_seedless_condensers] provided such a transformation only for $n\geq k\geq\Omega(\ell)$. Hence, our transformation makes a major improvement on their parameters. Such an improvement allows us to obtain better condensers for low-entropy oNOSF sources in the regime $n=\operatorname{poly}(\log(\ell/\varepsilon))$ (see 2).
1.3 Limits on Extraction from oNOSF Sources
Next we discuss our results on the limits of extraction from oNOSF sources . Our lower bounds are based on a new notion of influence of functions, namely online influence, that we introduce and analyze. Our upper bound results (explicit extractors) are based on a novel connection to leader election and coin-flipping protocols; to instantiate this connection and give explicit extractors, we construct the necessary protocols.
@@ 1.3 Limits on Extraction from oNOSF Sources
For simplicity, let’s focus on the case of $n=1$, which leads to interesting new questions about Boolean functions. We refer to such uniform ${\left(g,\ell,1\right)}$-oNOSF sources as ${\left(g,\ell\right)}$-oNOBF sources; oNOBF stands for online non-oblivious bit-fixing sources. We ask what is the exact tradeoff between $g$, $\ell$, and $\varepsilon$ for extracting from oNOBF sources . Towards this, we introduce the notion of online influence.
@@
We believe this is an interesting new measure and is worth studying in its own right. We refer the reader to Example 7.6 for a couple of interesting examples. For monotone functions (and more generally, unate functions), it is not hard to see that online influence equals the usual notion of influence (see Lemma 7.4 for a proof). Thus, to find interesting properties of online influence (compared to standard influence, Definition 7.1), one must look at non-monotone (in fact, non-unate) Boolean functions.
The following natural question arises towards our goal of proving extractor lower bounds: for a function $f$, what is the maximum online influence out of all $n$ bits? For the usual notion of influence, this question was resolved by the well-known theorem of [kahn_influence_1988], who showed there always exists a bit with influence at least $\operatorname*{Var}(f)\cdot\Omega\left(\frac{\log\ell}{\ell}\right)$.
We prove a Poincaré style inequality for total online influence, which shows that the above example is tight. For any function $f:\{0,1\}^{n}\rightarrow\{0,1\}$, define the function $e(f)(x)=(-1)^{f(x)}$.
Theorem 4 (Theorem 7.5, restated).
Using the above result, we obtain the following extraction lower bound.
@@
A similar extraction lower bound was shown in [aggarwal_how_2020] using different techniques.
@@
We now move on to presenting our explicit constructions of extractors for oNOBF and oNOSF sources. The following are our main results.
@@
There exists an explicit function $\mathsf{Ext}:\{0,1\}^{\ell}\to\{0,1\}$ such that for any ${\left(g,\ell\right)}$-oNOBF source $\mathbf{X}$ where $g\geq\ell-\ell/C\log(\ell)$, we have $\mathsf{Ext}(\mathbf{X})\approx_{\varepsilon=1/100}\mathbf{U}_{1}$, where $C$ is a large constant.
@@
There exists an explicit function $\mathsf{Ext}:(\{0,1\}^{n})^{\ell}\to\{0,1\}^{n}$ such that for any ${\left(g,\ell,n\right)}$-oNOSF source $\mathbf{X}$ where $g\geq\ell-\ell/C\log^{*}(\ell)$ and $n\geq\log(\ell)$, we have $\mathsf{Ext}(\mathbf{X})\approx_{\varepsilon=1/100}\mathbf{U}_{n}$, where $C$ is a large constant.
It is instructive to contrast our results with the non-online setting (where adversarial bits may depend on any good bit). The non-online versions of oNOSF sources and oNOBF sources are called NOSF sources and NOBF sources . For both these sources, the current best extractors require $g\geq\ell-\frac{\ell}{(\log\ell)^{2}}$, which is much more than what our extractors for 5 and 6 require.
We contrast the results for both settings in Tables 1 and 2. In these tables, we are providing known upper and lower bounds on the value of $b(\ell)$, defined as the maximum number of bad symbols for which extraction is still possible with a small constant error — so lower bounds correspond to best known constructions of such functions and upper bounds refers to best known limitation of such functions. We write “$O(\ell)$” to mean “$c\ell$ for some small universal constant $c<1$”.
To interpret our results in terms of (online) influence of coalitions, it will be useful to extend the definition of online influence to subsets of coordinates.
Definition 1.7.
For any function $f:\Sigma^{\ell}\to\{0,1\}$, and any $B\subset[{\ell}]$, where $B=\{i_{1}<i_{2}<\ldots<i_{k}\}$, define $\operatorname{\mathbf{oI}}_{B}(f)$ as follows: an online adversary $\mathcal{A}$ samples a distribution $\mathbf{X}$ in online manner. It starts by sampling the variables $x_{1},x_{2},\ldots,x_{i_{1}-1}$ independently and uniformly from $\Sigma$, then picking the value of $x_{i_{1}}$ depending on $x_{<i_{1}}$. Next, the variables $x_{i_{1}+1},\ldots,x_{i_{2}-1}$ are sampled independently and uniformly from $\Sigma$, and $\mathcal{A}$ sets the value of $x_{i_{2}}$ based on all set variables so for, and so on. Define the advantage of $\mathcal{A}$ to be $\textrm{adv}_{f,B}(\mathcal{A})=|\operatorname*{\mathbb{E}}[f(\mathbf{X})]-\operatorname*{\mathbb{E}}[f(\mathbf{U}_{\ell})]|$. Then $\operatorname{\mathbf{oI}}_{B}(f)$ is defined to be $\max_{\mathcal{A}}\{\textrm{adv}_{f,B}(\mathcal{A})\}$, where the maximum is taken over all online adversaries $\mathcal{A}$ that control the bits in $B$.
@@
NOBF $\Omega\left(\frac{\ell}{\log^{2}\ell}\right)$, [ajtai_influence_1993] $O\left(\frac{\ell}{\log\ell}\right)$, [kahn_influence_1988]
NOSF $\Omega\left(\frac{\ell}{\log^{2}\ell}\right)$, [ajtai_influence_1993] $O(\ell)$, [bourgain_influence_1992]
@@
oNOBF $\Omega\left(\frac{\ell}{\log\ell}\right)$, [Theorem 8.2] $O(\ell)$, Corollary 7.21 or [aggarwal_how_2020]
oNOSF $\Omega\left(\frac{\ell}{\log^{*}\ell}\right)$, [Theorem 8.3]55footnotemark: 5 $O(\ell)$, [aggarwal_how_2020]
@@
22footnotetext: Recall that this lower bound is for ${\left(g,\ell,n\right)}$-oNOSF sources with $n\geq\log(\ell)$.

### v3

@@ 1.2.2 Transforming Low-Entropy oNOSF sources to uniform oNOSF sources
Previously, [CGR24] provided such a transformation only for $n\geq k\geq\Omega(\ell)$. Hence, our transformation makes a major improvement on their parameters. Such an improvement allows us to obtain better condensers for low-entropy oNOSF sources in the regime $n=\operatorname{poly}(\log(\ell/\varepsilon))$ (see Theorem 4).
1.2.3 Existential Condensers
Theorem 3 (Informal version of Theorem 7.1).
For all $\ell,\varepsilon$ where $\ell\geq O(\log(1/\varepsilon))$, and $n=10^{4}$, there exists a condenser $\mathsf{Cond}:(\{0,1\}^{n})^{\ell}\to\{0,1\}^{m}$ such that for any uniform $\left(0.51\ell,\ell,n\right)$}-oNOSF source $\mathbf{X}$, we have $H_{\infty}^{\varepsilon}(\mathsf{Cond}(\mathbf{X}))\geq 0.99m$ where $m=\Omega(\ell+\log(1/\varepsilon))$. Furthermore, when $n=\omega(1)$, the output entropy rate becomes 1o(1)1-o(1).
This is tight since [CGR24] showed it is impossible to condense uniform $\left(0.5\ell,\ell,n\right)$}-oNOSF sources beyond output entropy rate $0.5$.
Theorem 4.
For all $\ell,n,\varepsilon$ where $n=\operatorname{poly}(\log(\ell/\varepsilon)),k=\Omega(\log(\ell/\varepsilon))$, there exists a condenser $\mathsf{Cond}:(\{0,1\}^{n})^{\ell}\to\{0,1\}^{m}$ such that for any $\left(0.51\ell,\ell,n,k\right)$}-oNOSF source $\mathbf{X}$, we have $H_{\infty}^{\varepsilon}(\mathsf{Cond}(\mathbf{X}))\geq m-O(m/\log(m))-O(\log(1/\varepsilon))$ where $m=\Omega(k)$.
We sketch the proof of both of these theorems in Section 2.1 We can also extend our result to condense from uniform $\left(g,\ell,n\right)$}-oNOSF sources for all $g,\ell$ and constant $n$ where the output entropy rate is $1/\left\lfloor\ell/g\right\rfloor-0.001$. This is tight since [CGR24] showed it is impossible to condense such sources beyond output entropy rate $1/\left\lfloor\ell/g\right\rfloor$.
Previously, [CGR24] showed how to existentially condense from uniform $\left(g,\ell,n\right)$}-oNOSF sources when $g\geq 0.51\ell$, provided $n\geq 2^{\omega(\ell)}$. As $n$ gets smaller, condensing becomes harder since a uniform $\left(g,\ell,n\right)$}-oNOSF source is also a uniform $\left(g\cdot n/1000,\ell\cdot n/1000,1000\right)$}-oNOSF source. Hence, we greatly improve the parameters while using different and much simpler techniques.
1.3 Extraction from oNOSF Sources
Next we discuss our positive and negative results on the limits of extraction from oNOSF sources . Our upper bound results (explicit extractors) are based on a novel connection to leader election and coin-flipping protocols; to instantiate this connection and give explicit extractors, we construct novel protocols for these distributed problems. Our lower bounds are based on a new notion of influence of functions, namely online influence, that we introduce and analyze.
@@ 1.3 Extraction from oNOSF Sources
For simplicity, we focus on the case of $n=1$, which leads to interesting new questions about Boolean functions. We refer to such uniform $\left(g,\ell,1\right)$}-oNOSF sources as $\left(g,\ell\right)$}-oNOBF sources; oNOBF stands for online non-oblivious bit-fixing sources. We ask what is the exact tradeoff between $g$, $\ell$, and $\varepsilon$ for extracting from oNOBF sources . Towards this, we introduce the notion of online influence.
@@
We establish new structural results on online influence, including a Poincaré-style inequality and use them to obtain the following extraction lower bound.
@@
A similar extraction lower bound was shown in [AORSV20] using different techniques.
@@
Here we present our explicit constructions of extractors for oNOBF and oNOSF sources. The following are our main results.
@@
There exists an explicit function $\mathsf{Ext}:\{0,1\}^{\ell}\to\{0,1\}$ such that for any $\left(g,\ell\right)$}-oNOBF source $\mathbf{X}$ where $g\geq\ell-\ell/(C\log(\ell))$, we have $\mathsf{Ext}(\mathbf{X})\approx_{\varepsilon=1/100}\mathbf{U}_{1}$, where $C$ is a large constant.
@@
There exists an explicit function $\mathsf{Ext}:(\{0,1\}^{n})^{\ell}\to\{0,1\}^{n}$ such that for any $\left(g,\ell,n\right)$}-oNOSF source $\mathbf{X}$ where $g\geq\ell-\ell/(C\log^{*}(\ell))$ and $n\geq\log(\ell)$, we have $\mathsf{Ext}(\mathbf{X})\approx_{\varepsilon=1/100}\mathbf{U}_{n}$, where $C$ is a large constant.
It is instructive to contrast our results with the non-online setting (where adversarial bits may depend on any good bit), called NOSF sources and NOBF sources . For both these sources, the current best extractors require $g\geq\ell-\frac{\ell}{(\log\ell)^{2}}$, which is much more than what Theorem 5 and Theorem 6 require.
We contrast the results for both settings in Tables 1 and 2. In these tables, we are providing known upper and lower bounds on the value of $b(\ell)$, defined as the maximum number of bad symbols for which extraction is still possible with a small constant error — so lower bounds correspond to best known constructions of such functions and upper bounds refers to the best known limitation of such functions. We write $O(\ell)$” to mean $c\ell$ for some small universal constant c<1c<1”.
To interpret our results in terms of (online) influence of coalitions, it will be useful to extend the definition of online influence to subsets of coordinates, which we do formally in Definition 8.1. Intuitively, we’re measuring the influence of the exact same adversary as in an oNOSF source .
@@
NOBF $\Omega\left(\frac{\ell}{\log^{2}\ell}\right)$, [AL93] $O\left(\frac{\ell}{\log\ell}\right)$, [KKL88]
NOSF $\Omega\left(\frac{\ell}{\log^{2}\ell}\right)$, [AL93] $O(\ell)$, [BKKKL92]
@@
oNOBF $\Omega\left(\frac{\ell}{\log\ell}\right)$, [Theorem 8.3] $O(\ell)$, Corollary 10.21 or [AORSV20]
oNOSF $\Omega\left(\frac{\ell}{\log^{*}\ell}\right)$, [Theorem 8.4] [^8] Recall that this lower bound is for $\left(g,\ell,n\right)$}-oNOSF sources with $n\geq\log(\ell)$. $O(\ell)$, [AORSV20]